# CLARIFID: Improving Radiology Report Generation by Reinforcing Clinically Accurate Impressions and Enforcing Detailed Findings

### v1

@@ Introduction
During training, we insert <next> tokens between consecutive findings sentences and prepend a special <impression> token before the impression section. This structured approach helps the model learn section-specific style and content, distinguishing between the detailed, observational language of Findings and the concise, conclusive language of Impressions. It also establishes clear section boundaries, preventing content bleeding between sections that could compromise clinical accuracy. Additionally, these special tokens create natural breakpoints for controlled decoding during inference, enabling our next-token forcing strategy.
@@
The decoder is trained autoregressively with a cross-entropy objective conditioned on the multi-image encoder’s outputs:
$\mathcal{L}_{\mathrm{CE}}=-\sum_{t=1}^{T}\log P_{\theta}\bigl{(}y_{t}\mid y_{<t},\widehat{\mathbf{S}},\mathbf{A}\bigr{)},$ (5)
where $y_{t}$ is the target token at step $t$, $\widehat{\mathbf{S}}$ and $\mathbf{A}$ are the padded image features and attention mask from the multi-image encoder, and $\theta$ represents the model parameters. This supervised pretraining stage establishes a foundation of radiological knowledge and reporting conventions before the reinforcement learning phase.
Impression-Based Reinforcement Learning
@@
After supervised pretraining, we employ Proximal Policy Optimization (PPO) to directly optimize the clinical utility of generated reports. Unlike previous approaches that rely on text-similarity metrics or entity-level rewards, we specifically target the Impression section’s clinical accuracy using CheXbert-based F1 scores as the primary reward signal.
PPO Framework
@@
PPO (Schulman et al. 2017) is a policy gradient method that provides stable updates by constraining the policy change between iterations. For our report generation task, we conceptualize the problem in terms of states, actions, policy, and rewards. The states ($s_{t}$) consist of the current partial report and visual features. Actions ($a_{t}$) represent the selection of the next token in the sequence. The policy ($\pi_{\theta}$) defines the probability distribution over possible next tokens. The reward ($r$) is derived from the CheXbert-based F1 score calculated on the Impression section.
The PPO objective is formulated as:
@@
$\mathcal{L}_{PPO}=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\,\text{clip}\left(r_{t}(\theta),\,1-\epsilon,\,1+\epsilon\right)\hat{A}_{t}\right)\right]$ (6)
where the probability ratio $r_{t}(\theta)$ between the current and old policies is defined as:
$r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ (7)
Reward Formulation
@@
After generating a complete report, we extract the Impression section and evaluate it using the CheXbert model (Smit et al. 2020), which assigns labels to 14 common radiological conditions. The reward is calculated as the micro-averaged F1 score between CheXbert-extracted labels from the generated impression and the ground-truth labels:
$r=\text{F1}(\text{CheXbert}(\text{Impression}_{gen}),\text{Labels}_{gt})$
By optimizing for impression accuracy rather than token-level likelihood or findings fidelity, we directly align the model with clinical decision-making, implicitly encouraging the generation of findings that support accurate diagnostic conclusions.
@@
0:  generator $A$, value model $V$
1:  parameters: $k$ required <next> tokens, $N$ candidates, $T_{\text{find}}$, $T_{\text{imp}}$
2:  for $j\leftarrow 1$ to $N$ do
3:     $\mathbf{t}\leftarrow$Findings:
@@
5:        $y\leftarrow A.\text{sample\_next}(\mathbf{t},T_{\text{find}})$
[…]
8:        end if
9:        $\mathbf{t}\leftarrow\mathbf{t}\,\|\,y$
10:     end while
[…]
12:     repeat
13:        $y\leftarrow A.\text{sample\_next}(\mathbf{t},T_{\text{imp}})$
14:        $\mathbf{t}\leftarrow\mathbf{t}\,\|\,y$
[…]
16:     $\hat{s}_{j}\leftarrow\mathbf{t}$,    $v_{j}\leftarrow V(\hat{s}_{j})[-1]$
17:  end for
18:  $j^{\ast}\leftarrow\arg\max_{j}\,v_{j}$
19:  return  $\hat{s}\leftarrow\hat{s}_{j^{\ast}}$
Inference-Time Control
Next-Token Forcing
To ensure sufficiently detailed reports, we adopt a controlled decoding scheme using <next> and <impression> tokens. Impression-related tokens (including <impression> and <eos>) are blocked until the Findings section exceeds a minimum length, preventing early termination and yielding richer, multi-sentence findings. Motivated by budget forcing (Muennighoff et al. 2025), we treat the Findings section as intermediate reasoning and the Impression section as the final answer, encouraging grounded and interpretable outputs.
Best-of-N Decoding with Temperature Scheduling
We generate the full Findings–Impression report in a single pass, but apply two different sampling temperatures:
  • Findings. A slightly higher temperature $T_{\text{find}}$ encourages diversity and richer detail, helping the model surface subtle abnormalities.
  • Impression. A lower temperature $T_{\text{imp}}$ promotes consistency in the clinical summary.
@@
Table 1: Comparison with SOTA methods on MIMIC-CXR (M-CXR), MIMIC-CXR Impression (M-CXR Imp.), and IU X-ray. Methods annotated with the symbol †have been re-evaluated using official codes. The best performance for each dataset is highlighted in bold.
@@
Table 2: Results of ablation experiments on MIMIC-CXR Imp. dataset. The best results are highlighted in bold.
Experiments
Dataset
MIMIC-CXR
We use the MIMIC-CXR (Johnson et al. 2019) dataset, which contains a large collection of chest X-ray images, along with corresponding radiology reports. MIMIC-CXR groups X-ray images and their associated radiology reports by “study.” Each study may contain multiple chest X-ray images, along with a single radiology report. We follow the standard train, validation, and test splits provided in (Johnson et al. 2019). For fair comparison with prior works, our evaluation on the full test set compares only the generated ”Findings” section with the ground-truth ”Findings” section.
MIMIC-CXR Imp.
Under official splitting of MIMIC-CXR, we specifically select only those reports containing both the ”Findings” and ”Impression” sections, ensuring that our model is trained on comprehensive reports reflective of real-world clinical workflows.
IU X-ray
To assess the ability to generalize of our approach, we further evaluate it on the public IU X-ray collection (Demner-Fushman et al. 2016). Following the protocol proposed by (Jin et al. 2024), we train the model exclusively on MIMIC-CXR and evaluate it on the IU X-ray test split defined in that work. For a fair comparison with MIMIC-CXR, we keep only those studies that include both a “Findings” and an “Impression” section, resulting in 4,162 chest-radiograph images paired with 2,084 reports.
Figure 4: Section-wise F1 scores on the MIMIC-CXR Impression test set through RL iteration.
Evaluation Metrics
Following previous works (Chen et al. 2022b, a; Huang, Zhang, and Zhang 2023; Gu et al. 2024), we measure both clinical efficacy (CE) and natural language generation (NLG) metrics. Specifically, we compute Precision, Recall, and F1-scores (P, R, F1) for CE, and BLEU-1(B-1), BLEU-4(B-4), METEOR(MTR), and ROUGE-L(R-L) for NLG. For CE, we utilize CheXbert to extract disease labels from the generated text and then calculate micro-averaged P, R, and F1 following (Chen et al. 2022b, a). NLG metrics are computed using the (Lin et al. 2015). In earlier studies that focus only on generating “Findings” (without an “Impression”), evaluation is performed solely on the “Findings” section. For models that also produce “Impression”, we consider both “Findings” and “Impression”.
Implementation details
All components used in our experiments are trained from scratch. We use GPT-2 small (Radford et al. 2019) as the text-decoder backbone and SwinV2 (Liu et al. 2022) as the image-encoder backbone. Throughout both training and inference, all images are resized to $H{=}512,\,W{=}512$. In PPO training phase we use $\epsilon=0.2$ and a KL-divergence penalty coefficient of 0.05. At inference, the number of forced <next> tokens was 10, $N$=8 for best-of-N, and $T_{\text{find}}=1.0,T_{\text{imp}}=0.8,p=0.9$ for temperature scheduling. Further details are provided in Appendix.
Result
Comparison with SOTA

### v4

@@ 3.2 Section-aware pretraining
After image encoder warm-up, to explicitly represent the two-stage structure of radiology reports, consisting of detailed observations (Findings) followed by a diagnostic summary (Impression), we then adopt a section-aware pretraining scheme that uses control tokens to mark the report’s logical structure at the sentence level. During training, we insert <next> tokens between consecutive findings sentences and prepend a special <impression> token before the impression section. This structured approach helps the model learn section-specific style and content, distinguishing between the detailed, observational language of Findings and the concise, conclusive language of Impressions. It also establishes clear section boundaries, preventing content bleeding between sections that could compromise clinical accuracy. Additionally, these special tokens create natural breakpoints for controlled decoding during inference, enabling our next-token forcing strategy.
@@
The decoder is trained with a cross-entropy objective, jointly optimized with the classification loss $\mathcal{L}_{\mathrm{cls}}$, conditioned on the multi-image encoder’s outputs:
$\mathcal{L}_{\mathrm{CE}}$ $=-\sum_{t=1}^{T}\log P_{\theta}\bigl(y_{t}\mid y_{<t},\widehat{\mathbf{S}},\mathbf{A}\bigr)$ (7)
$\mathcal{L}_{\mathrm{pre}}$ $=\mathcal{L}_{\mathrm{cls}}+\mathcal{L}_{\mathrm{CE}}$ (8)
Where $\mathcal{L}_{\mathrm{CE}}$ is the cross-entropy loss, $\mathcal{L}_{\mathrm{cls}}$ is the weighted binary cross-entropy loss for image classification, and $\mathcal{L}_{\mathrm{pre}}$ is the joint objective for section-aware pretraining. This supervised pretraining stage establishes a foundation of radiological knowledge and reporting conventions before the reinforcement learning phase.
Algorithm 1 CLARIFID Training
0: Dataset $\mathcal{D}$, generator $\pi_{\theta}$, generator image encoder $E_{\theta}$, value model $V_{\phi}$, pretraining cycles $C_{\mathrm{pre}}$, image-only cycles $C_{\text{img}}$, PPO iterations $C_{\mathrm{ppo}}$, Value function loss weight $\lambda_{V}$
1:Stage I: Section-aware supervised pretraining
2:Init: Add the special tokens <next> and <impression> to the tokenizer.
3:for $\mathrm{iter}\leftarrow 1$ to $C_{\mathrm{pre}}$ do
4:  Repeat over mini-batches.
5:  Encode images to get $(\widehat{\mathbf{S}},\mathbf{A})$.
6:  Compute 14 logits using (4).
7:  Binarize ground-truth label with (5)
8:  if $\mathrm{iter}\leq C_{\text{img}}$ then
9:   # image encoder warm-up
10:   Update $E_{\theta}$ using the weighted BCE loss in (6).
11:  else
12:   # section-aware LM pretraining
13:   Build targets: insert <next> in Findings, prepend <impression>.
14:   Update $\pi_{\theta}$ with joint objective (8) conditioned on $(\widehat{\mathbf{S}},\mathbf{A})$.
15:  end if
16:end for
17:Stage II: Impression-based reinforcement learning
[…]
19:for $\mathrm{iter}\leftarrow 1$ to $C_{\mathrm{ppo}}$ do
20:  Collect a set of trajectories $\mathcal{T}$ from interactions using current policy $\pi_{\theta}$
21:  for each trajectory $\tau\in\mathcal{T}$ do
22:   $r\leftarrow\mathrm{F1}\!\big(\mathrm{CheXbert}(\mathrm{Impression}_{\text{gen}}),\mathrm{CheXbert}(\mathrm{Impression}_{\text{gt}})\big)$
23:   Compute augmented rewards $R^{\prime}$ using  (9)
24:   Estimate advantages $\hat{A}$ with (10) – (11)
25:  end for
26:  for $\mathrm{epoch}\leftarrow 1$ to $E_{\mathrm{ppo}}$ do
27:   for each mini-batch of data from $\mathcal{T}$ do
28:    Update $\theta,\phi$ by optimizing $\mathcal{L}_{\text{PPO}}+\lambda_{V}\mathcal{L}_{{V}}$ from (12) – (14)
29:   end for
30:  end for
31:end for
32:return $\pi_{\theta}$
3.3 Impression-based reinforcement learning
@@
After supervised pretraining, we employ Proximal Policy Optimization (PPO) to directly optimize the clinical utility of generated reports. Unlike previous approaches that rely on text-similarity metrics or entity-level rewards, we specifically target the Impression section’s clinical accuracy using CheXbert-based F1 scores as the primary reward signal. We expect that optimizing Impression-level clinical accuracy as the reinforcement signal will encourage the model to refine the Findings section as an intermediate reasoning step, thereby improving its factual alignment with the final diagnostic conclusions.
3.3.1 PPO framework
@@
PPO Schulman et al. (2017) is a policy gradient method that provides stable updates by constraining the policy change between iterations. For our report generation task, we conceptualize the problem in terms of states, actions, policies, and rewards. The states ($s_{t}$) consist of the current partial report and visual features, while actions ($a_{t}$) represent the selection of the next token in the sequence. The policy ($\pi_{\theta}$) defines the probability distribution over possible next tokens.
Rollout and reward structure.
At each PPO iteration, we perform $N$ stochastic rollouts per input using temperature sampling, which increases on-policy exploration and provides a richer set of trajectories without additional gradient steps. The reward formulation is token-level, consisting of (i) a per-token KL penalty against a reference policy $\pi_{\text{ref}}$ to preserve fluency and (ii) a terminal clinical reward from the CheXbert-based F1 score on the Impression section.
Policy regularization.
A major challenge in reinforcement learning fine-tuning is preventing policy collapse, where the model sacrifices linguistic quality and coherence to maximize a sparse, terminal reward. To address this, we introduce KL-divergence regularization, which provides a dense reward signal to guide the policy at every generation step. This per-token KL term acts as a shaping reward, penalizing the current policy $\pi_{\theta}$ if it strays too far from the fluent distribution of the supervised pre-trained reference policy $\pi_{\text{ref}}$. The final reward is augmented as follows:
$R^{\prime}=R-\beta\cdot\mathrm{KL}(\pi_{\theta}\,\|\,\pi_{\text{ref}})$ (9)
Here, $\beta$ is a fixed coefficient that balances the trade-off between maximizing the clinical accuracy reward ($R$) and maintaining the linguistic quality inherited from the reference model. This provides a stable regularization effect throughout the fine-tuning process.
Generalized advantage estimation.
A key component of PPO training is the estimation of the advantage function, which quantifies how much better (or worse) an action $a_{t}$ is compared to the expected value of state $s_{t}$. Directly using Monte Carlo returns provides unbiased estimates but suffers from very high variance, while bootstrap-based temporal-difference methods exhibit low variance but are biased. To balance this trade-off, we adopt Generalized Advantage Estimation (GAE) Schulman et al. (2015), which interpolates between these two extremes using the parameter $\lambda\in[0,1]$. To estimate advantages, we use Generalized Advantage Estimation (GAE) with discount factor $\gamma=1$ and trace-decay parameter $\lambda=0.95$:
$\delta_{t}$ $=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),\qquad\hat{A}_{t}=\sum_{l=0}^{T-t-1}(\gamma\lambda)^{l}\delta_{t+l}$ (10)
We then apply advantage whitening Schulman et al. (2017); Dhariwal et al. (2017); Huang et al. (2024) over non-padded positions:
$\tilde{A}_{t}=\frac{\hat{A}_{t}-\mu_{A}}{\sigma_{A}}$ (11)
which reduces variance and improves numerical stability.
Objective function.
Finally, the policy parameters are updated by maximizing the clipped PPO Schulman et al. (2017) surrogate objective. The standard policy gradient objective can be written in terms of the probability ratio $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$, which measures how much the new policy deviates from the old policy for the same action. If $r_{t}(\theta)>1$, the action becomes more likely under the new policy, and if $r_{t}(\theta)<1$, it becomes less likely.
Directly optimizing $r_{t}(\theta)\tilde{A}_{t}$ can lead to excessively large updates when $r_{t}(\theta)$ deviates far from 1, causing training instability. To mitigate this, PPO introduces a clipping mechanism:
@@
$\mathcal{L}_{\mathrm{PPO}}=\mathbb{E}_{t}\!\left[\min\Big(r_{t}(\theta)\,\tilde{A}_{t},\;\operatorname{clip}\!\big(r_{t}(\theta),\,1-\epsilon,\,1+\epsilon\big)\,\tilde{A}_{t}\Big)\right]$ (12)
where $\epsilon$ limits how much the new policy is allowed to diverge from the old one. The clipping ensures that when $r_{t}(\theta)$ goes outside the range $1-\epsilon,1+\epsilon]$, the update is truncated, preventing the policy from moving too far in a single step.
Intuitively, this surrogate objective preserves the benefits of large advantage updates when the policy is close to the old one, while penalizing overly aggressive updates that might destabilize learning. Together with GAE and KL regularization, the clipped objective forms the core of PPO’s stability, balancing exploration with conservative policy improvement.
Value function training.
To ensure stable advantage estimates, the value function $V_{\phi}$ is trained alongside the policy. It is updated by minimizing a clipped mean squared error loss, which prevents the value network from undergoing drastic changes that could destabilize policy updates. The target for this update is empirical return $R_{t}=\hat{A}_{t}+V_{\mathrm{old}}(s_{t})$, where $V_{\mathrm{old}}(s_{t})$ represents the value estimate from before the update step. Similar to the policy objective, the new value prediction $V_{\phi}(S_{t})$ is clipped to stay within an $\epsilon$-ball of the old value $V_{\phi_{old}}(S_{t})$.
$\mathcal{L}_{{V}}=\frac{1}{2}\,\mathbb{E}_{t\in\mathcal{M}}\Big[\max\!\Big(\big(V_{\phi}(s_{t})-R_{t}\big)^{2},\;\big(\mathrm{clip}(V_{\phi}(s_{t}),\,V_{\mathrm{old}}(s_{t}){\pm}\epsilon\big)-R_{t})^{2}\Big)\Big]$ (13)
At each update, we optimize a joint objective combining the clipped PPO surrogate and the value loss:
$\mathcal{L}_{\text{total}}\;=\;\mathcal{L}_{\text{PPO}}\;+\;\lambda_{V}\,\mathcal{L}_{{V}}$ (14)
Where $\lambda_{V}$ is the value loss weight.
3.3.2 Reward formulation
@@
After generating a complete report, we extract the Impression section and evaluate it using the CheXbert model, which assigns labels to 14 common radiological conditions. The reward is calculated as the micro-averaged F1 score between the CheXbert-extracted labels from the generated impression and the ground truth. To stabilize optimization, we apply advantage whitening after estimating returns, reducing variance across trajectories while preserving the natural scale of the CheXbert-F1 reward.
$r=\text{F1}(\text{CheXbert}(\text{Impression}_{gen}),\text{CheXbert}(\text{Impression}_{gt}))$
Algorithm 1 summarizes the entire training pipeline, from section-aware pretraining to impression-based reinforcement learning.
@@
0: Generator $\pi_{\theta}$, value model $V_{\phi}$, required <next> token count $k$, number of candidates NN, temperatures $T_{\mathrm{find}},T_{\mathrm{imp}}$
1:for $j\leftarrow 1$ to NN do
2:  $\mathbf{t}\leftarrowFindings$:
3:  while $\texttt{<next>}(\mathbf{t})<k$ do
4:   $y\leftarrow\pi_{\theta}.\text{sample\_next}(\mathbf{t},T_{\mathrm{find}})$
5:   if $y\in\{\texttt{<eos>},\texttt{<impression>}\}$ then
6:    $y\leftarrow\texttt{<next>}$
7:   end if
8:   Append $y$ to $\mathbf{t}$
9:  end while
10:  Append <impression> to $\mathbf{t}$
11:  repeat
12:   $y\leftarrow\pi_{\theta}.\text{sample\_next}(\mathbf{t},T_{\mathrm{imp}})$
13:   Append $y$ to $\mathbf{t}$
14:  until $y=\texttt{<eos>}$
15:  $\hat{s}_{j}\leftarrow\mathbf{t}$,  $v_{j}\leftarrow V_{\phi}(\hat{s}_{j})[-1]$
16:end for
@@
18:return $\hat{s}_{j^{\ast}}$
3.4 Inference-time control
3.4.1 Next-token forcing
To ensure sufficiently detailed reports, we adopt a controlled decoding scheme using <next> and <impression> tokens. Impression-related tokens (including <impression> and <eos>) are blocked until the Findings section exceeds a minimum length, preventing early termination and yielding richer, multi-sentence findings. Motivated by budget forcing Muennighoff et al. (2025), we treat the Findings section as intermediate reasoning and the Impression section as the final answer, encouraging grounded and interpretable outputs.
3.4.2 Best-of-N decoding with temperature scheduling
We generate the full Findings–Impression report in a single pass while sampling with distinct temperatures for each section. For the Findings, we use a slightly higher temperature $T_{\text{find}}$ to encourage diversity and richer detail, helping the model surface subtle abnormalities. For the Impression, we use a lower temperature $T_{\text{imp}}$ to promote consistency in the clinical summary.
@@ 3.4.2 Best-of-N decoding with temperature scheduling
4 Experiments
4.1 Dataset
4.1.1 MIMIC-CXR
We use the MIMIC-CXR Johnson et al. (2019) dataset, which contains a large collection of chest X-ray images, along with corresponding radiology reports. MIMIC-CXR groups X-ray images and their associated radiology reports by “study.” Each study may contain multiple chest X-ray images, along with a single radiology report. We follow the standard train, validation, and test splits provided in Johnson et al. (2019). For fair comparison with prior works, our evaluation on the full test set compares only the generated “Findings” section with the ground-truth “Findings” section.
4.1.2 MIMIC-CXR Impression
Under official splitting of MIMIC-CXR, we specifically select only those reports containing both the “Findings” and “Impression” sections, ensuring that our model is trained on comprehensive reports reflective of real-world clinical workflows. After filtering, we obtain 123,742 training, 977 validation, and 1,600 test studies.
4.1.3 IU X-ray
To assess the generalization ability of our approach, we further evaluate it on the public IU X-ray collection Demner-Fushman et al. (2015). Following the protocol proposed by Jin et al. (2024), we train the model exclusively on MIMIC-CXR and evaluate it on the IU X-ray test split defined in that work. For a fair comparison with MIMIC-CXR, we keep only those studies that include both a “Findings” and an “Impression” section, resulting in 4,162 chest-radiograph images paired with 2,081 reports.
4.2 Evaluation metrics
Following previous works Chen et al. (2020, 2022); Huang et al. (2023); Gu et al. (2023), we measure both clinical efficacy (CE) and natural language generation (NLG) metrics. Specifically, we compute Precision, Recall, and F1-scores (P, R, F1) for CE, and BLEU-1(B-1), BLEU-4(B-4), METEOR(MTR), and ROUGE-L(R-L) for NLG. For CE, we utilize CheXbert to extract disease labels from the generated text and then calculate micro-averaged P, R, and F1 following Chen et al. (2020, 2022). NLG metrics are computed using the Lin et al. (2014). In earlier studies that focus only on generating “Findings” (without an “Impression”), evaluation is performed solely on the “Findings” section. For models that also produce “Impression”, we consider both “Findings” and “Impression”.
4.3 Implementation details
All components are trained from scratch. We use the AdamW optimizer Loshchilov and Hutter (2017). All images are resized to $512\times 512$ during training and inference. During section-aware pretraining we optimize the image encoder, the lightweight aggregator, and the GPT-2 decoder end-to-end with a learning rate of $1\times 10^{-4}$. To improve robustness we apply light geometric augmentation with random rotation within $\pm 5^{\circ}$. At each epoch we randomly permute the order of Findings sentences, insert <next> between adjacent sentences, and prepend <impression> before the Impression. The policy model uses a 30-layer GPT-2 backbone and the value model uses a 6-layer backbone. We use $C_{\text{pre}}=68$ and $C_{\text{img}}=10$.
In the PPO phase we use a learning rate of $1\times 10^{-5}$, a KL divergence penalty coefficient of $0.05$, 4 rollouts per sample, $\epsilon=0.2$, and $\lambda_{V}=0.1$. The PPO batch size is 4 with gradient accumulation for an effective batch size of 256. Sampling temperature during rollouts is fixed at 1.0.
At inference we use best-of-N with $N=8$. We force up to 10 <next> tokens to maintain sentence boundaries in Findings. Temperatures are $T_{\text{find}}=1.0$ and $T_{\text{imp}}=0.8$ with nucleus sampling $p=0.9$.
4.3.1 Inference time
The evaluation was carried out on a single NVIDIA RTX 6000 GPU with a batch size of 2. On average per study, processing the MIMIC-CXR Impression test set took 0.90 s with plain greedy decoding, 6.43 s with next-token forcing to enforce the Findings length, and 13.41 s with both next-token forcing and the BoN strategy.
Table 1: Comparison with SOTA methods on MIMIC-CXR (M-CXR), MIMIC-CXR Impression (M-CXR Imp.), and IU X-ray. Methods annotated with the symbol $\dagger$ have been re-evaluated using official codes. The best performance for each dataset is in bold.
@@ 4.3.1 Inference time
Table 2: Results of ablation experiments on MIMIC-CXR Imp. dataset. The best results are in bold.
@@
5 Results
5.1 Comparison with SOTA